OpenStudy (anonymous):

-5<2m≤9 m is an integer. Write down all the possible values of m. This is my first time using this site. Please could someone explain what an integer is and how I can solve this problem. :)

OpenStudy (anonymous):

\[-5<2m \le9\] divide each inequality by 2 \[-\frac{ 5 }{ 2 }<m \le \frac{ 9 }{ 2 }\] \[-2.5<m \le 4.5\] m is an integer. \[m \in [-2,-1,0,1,2,3,4] \]

OpenStudy (butterflydreamer):

A simple definition of an integer is that it is a "WHOLE" number :) So anything in a fraction form, decimal form etc. So number like: